OpenStudy (anonymous):

Help.. The table shows the height of a plant as it grows. What equation in point-slope form gives the plant’s height at any time? Let y stand for the height of the plant in cm and let x stand for the time in months. Time (months) Plant Height (cm) 3 15 5 25 7 35 9 45

OpenStudy (anonymous):

y-15 = (25-15)/(5-3) (x-3) so y = 5x

OpenStudy (anonymous):

first make a graph... plot (3,15) , ( 5,25) , ( 7,35) and (9,45) points